A single slit diffraction pattern shown above. The light used to make the pattern has a wavelength of 461 nm and the screen is 1.2 m from the slit. If the slit has a width 9 mm, what is the width w of the central maximum in cm?
Answer #1

d*sinθ = m*λ

both sides of the screen we know that are minima and in between them is a central maximum

so first we calculate the distance of minima one side and then we double it s we can find the width of maximum laying   between them

GOT IT ?

so for minimum  

Y = m*λ*D/d = 1*461*10-9 *1.2/( 9*10^-3) =

= 6.1466*10-5 m

width w of the central maximum = 2*Y =1.229*10-4 m = 0.0122933 cm answer

****************************************************************************
Goodluck Comment in case any doubt, will reply for sure..